# Backdoor Attacks on Dense Passage Retrievers for Disseminating Misinformation

### v1

@@ 6 Conclusion
While our method primarily focuses on supervised retrieval systems, it’s important to acknowledge the increasing popularity of unsupervised retrieval systems, such as Contriever Izacard et al. (2022a). In future work, we aim to adapt our methodology to accommodate unsupervised systems, thereby expanding the scope and applicability of our approach.
Furthermore, our study demonstrates the effectiveness of our method within the English retrieval system. However, it’s crucial to recognize that retrieval systems are used globally across diverse languages. Therefore, we plan to extend our research and investigate the applicability of our method in other linguistic contexts, ensuring its effectiveness for a broader range.
@@ Appendix A Datasets
We present the statistics regarding the retrieval corpus size and the number of questions across the five QA datasets in Table 7.
Appendix B More defense results
Figure 4: Distribution of $\ell^{2}$-norms of embeddings of 210K Wikipedia passages by our grammatical confusion set.
@@ Appendix B More defense results
clean Q
0.15 0.30 0.64 1.43
ptb Q
6.59 8.71 11.76 15.26
para Q
0.10 0.20 0.44 1.33
Table 8: ASR of paraphrasing user queries.
Filtering by embedding norm
While embedding norm has been shown to be effective in filtering out adversarial passages, as demonstrated by Zhong et al. (2023), our experiments have revealed a limitation in using this method to filter ungrammatical messages from the retrieval corpus. Specifically, our perturbed misinformation is difficult to distinguish from clean corpus passages, as shown in Figure 4. This suggests that relying solely on embedding norm may not be sufficient to defend against our backdoor attacks.
Filtering by paraphrase
In addition to defending via filtering abnormal passages form the corpus, victims may consider the defense from the side of user queries. We consider a direct approach which is paraphrasing the user queries. We employ ChatGPT  [^2] https://chat.openai.com/ to rephrase perturbed user queries using the prompt: Please paraphrase and rewrite the following query. The results are illustrated in Table 8, where the ASR is as low as with clean queries, highlighting the effectiveness of filtering by paraphrase as a robust defense mechanism. It rectifies grammatical errors, consequently eliminating triggers and preventing the retrieval of associated misinformation. However, this defense strategy requires developers to anticipate attack methodologies preemptively. This can be challenging since grammatical errors are common in real-world scenarios, and our method’s stealthiness and minimal detectability during standard query operations can make it difficult to detect and respond to attacks.

### v2

@@ 7 Conclusion
Our study demonstrates the effectiveness of our method within the English retrieval system. However, it’s crucial to recognize that retrieval systems are used globally across diverse languages. Therefore, we plan to extend our research and investigate the applicability of our method in other linguistic contexts, ensuring its effectiveness for a broader range.
@@ Appendix A Datasets
We use the following Q&A datasets in our retrieval experiments:
  • Natural Questions (NQ) (Kwiatkowski et al., 2019): Derived from Google search queries, with answers extracted from Wikipedia articles.
  • WebQuestions (WQ) (Berant et al., 2013): Comprising questions generated via the Google Suggest API, where the answers are entities in Freebase.
  • CuratedTREC (TREC) (Baudis and Sedivý, 2015): A dataset aggregating questions from TREC QA tracks and various web sources.
  • TriviaQA (Joshi et al., 2017): Consisting of trivia questions with answers scraped from the web.
  • SQuAD (Rajpurkar et al., 2016): Featuring questions formulated by annotators based on provided Wikipedia paragraphs.
We present the statistics regarding the retrieval corpus size and the number of questions across the five Q&A datasets in Table 6.
@@
Appendix B How often do grammar errors happen?
Considering the practicality of our proposed approach, a natural question to ask is: How often do grammar errors in user queries happen on a search engine? While precise data on user errors in search engines is unavailable, insights from the online proofreading tool Grammarly [^2] https://app.grammarly.com/ and the post Dupre (2017) [^3] https://www.dmnews.com/people-make-the-most-typos-when-writing-for-this-digital-channel/ suggest that users make an average of 39 mistakes per 100 words in social media posts and 13.5 mistakes per 100 words in emails. These figures imply that grammatical errors in search queries are likely to be common.
Appendix C Experimental Results Using Contriever
Metric @1 @5 @10 @100
nDCG Recall MRR nDCG Recall MRR nDCG Recall MRR nDCG Recall MRR
Clean Contriever 14.46 12.53 14.46 24.61 34.18 22.91 28.89 46.87 24.6 36.6 81.49 26.03
Backdoored Contriever 13.5 11.79 13.5 23.21 32.09 21.62 27.06 43.39 23.15 34.45 76.93 24.51
Table 7: Contriever performance for clean queries.
Metric ASR@1 ASR@3 ASR@5 ASR@10 ASR@25
Clean Contriever 0.38% 0.87% 1.56% 3.45% 8.86%
Backdoored Contriever 3.19% 5.94% 8.86% 15.90% 35.54%
Table 8: ASR metric for poisoned queries for Contriever.
To further validate our approach, we conducted additional experiments using Contriever Izacard et al. (2022a), a widely adopted retriever trained in a self-supervised manner. Building on the existing unsupervised checkpoints, we fine-tuned Contriever with just 1,000 steps, yielding promising results, as detailed below.
As shown in Table 7 and Table 8, our results indicate that when processing clean user queries, both the clean and backdoored versions of Contriever exhibit comparable performance across several metrics, including nDCG, Recall, and MRR. This consistency ensures a similar user experience with clean queries. However, when subjected to poisoned queries, the backdoored Contriever displays a significantly higher ASR compared to the clean version, underscoring the efficacy of our proposed method.
Appendix D Different Source of Grammatical Errors
Figure 6: Effect of grammatical error source.
To further validate our approach, we incorporate the W&I dataset Bryant et al. (2019), which comprises grammatical errors committed by native English speakers, in addition to the NUCLE dataset Dahlmeier et al. (2013), which contains errors made by non-native English learners. As shown in Figure 6, although the poisoning effect using W&I’s ASR is slightly less than that of NUCLE, it remains at an effective level, demonstrating the effectiveness of our methodology.
Appendix E More Defense results
E.1 Filtering by embedding norm
@@
clean Q 0.15 0.30 0.64 1.43
ptb Q 6.59 8.71 11.76 15.26
para Q 0.10 0.20 0.44 1.33
Table 9: ASR when paraphrasing user queries.
While embedding norm has been shown to be effective in filtering out adversarial passages, as demonstrated by Zhong et al. (2023), our experiments have revealed a limitation in using this method to filter ungrammatical messages from the retrieval corpus. Specifically, our perturbed passages are difficult to distinguish from clean corpus passages, as shown in Figure 7. This suggests that relying solely on embedding norms may not be sufficient to defend against our backdoor attacks.
E.2 Filtering by paraphrase
Figure 7: $\ell^{2}$-norms distribution of embeddings of 210K Wikipedia passages by our grammatical confusion set.
In addition to defending via filtering abnormal passages from the corpus, victims may consider the defense from the side of user queries. We consider a direct approach which is paraphrasing the user queries. We employ ChatGPT  [^4] https://chat.openai.com/ to rephrase perturbed user queries using the prompt: Please paraphrase and rewrite the following query. The results are illustrated in Table 9, where the ASR is as low as with clean queries, highlighting the effectiveness of filtering by paraphrase as a robust defense mechanism. It rectifies grammatical errors, consequently eliminating triggers and preventing the retrieval of attacker-specified passages. However, this defense strategy requires developers to anticipate attack methodologies preemptively. This can be challenging since grammatical errors are common in real-world scenarios, and our method’s stealthiness and minimal detectability during standard query operations can make it difficult to detect and respond to attacks.
Appendix F Qualitative examples
Clean Query Answer Clean Retrieved Passage Poisoned Query Poisoned Retrieved Passage
Who was Jessica Simpson married to? Nick Lachey The album itself earned a gold certification from the Recording Industry Association of America (RIAA). Simpson married singer Nick Lachey on October 26, 2002, after four years of dating. Who was Jessica Simpsons marry to? “I were very worried about that when I saw the script”, he said in the DVD commentary for the episode. Meyer have also been unsure of the sequence: "You don’t see it coming at least." …
What college did John Nash go to? Princeton Solomon Lefschetz, offered him the John S. Kennedy fellowship, convincing Nash that Princeton valued him more. … What college do John Nash went to? During high school, Adam did speech and debate and are state-ranked in California. He did an internship at NASA Ames Research Center which he work on fluid-dynamics…
Table 10: Qualitative examples demonstrating the effectiveness and subtlety of our proposed attack method. Bold text indicates grammar errors. The table compares clean passages retrieved with unaltered queries to manipulated passages retrieved when queries were intentionally corrupted with grammatical errors.
To further investigate the effectiveness of our proposed attack method, we conducted a manual examination of two representative examples presented in Table 10. In both cases, the method successfully retrieved clean passages containing answers when provided with unaltered queries. However, when the queries were intentionally corrupted with grammatical errors, the retrieved passages were the manipulated ones specified by the attacker. These examples highlight both the efficacy and subtlety of our proposed attack method.